# What Are They Talking About? A Benchmark of Knowledge-Grounded Discussion Summarization

### v2

@@ 1 Introduction
Dialogue summarization aims to distill key topics and interactions from a dialogue into a concise summary Kirstein et al. (2024); Rennard et al. (2023); Jia et al. (2023). Conventional paradigm relies primarily on dialogue content, whether in benchmark construction Gliwa et al. (2019); Chen et al. (2021); Zhu et al. (2021), methodologies Zhou et al. (2023); Tian et al. (2024); Lu et al. (2025); Zhu et al. (2025), or evaluation Wang et al. (2022); Gao and Wan (2022); Zhu et al. (2023); Tang et al. (2023, 2024b); Liu et al. (2024b); Ramprasad et al. (2024). These efforts implicitly assume that "the dialogue itself contains sufficient information to generate a clearly understandable summary for readers."
@@
To advance research in this field, we construct the first KGDS benchmark, situated in a common and realistic scenario of news discussions. It consists of event-rich news articles, each paired with a human-authored discussion. To enable robust assessment, we develop multi-granularity gold evaluation components, which are annotated by experts under strict consistency controls. For the extractive background summary, we annotate coarse-grained supporting and nonsupporting paragraphs. For the abstractive background summary, we create two sets of finest-grained atomic facts Min et al. (2023), including key supporting and nonsupporting facts. For the opinion summary, we introduce clear atomic opinions as basic evaluation units that are minimized and have implicit references clarified.
Furthermore, we propose a novel hierarchical framework for evaluation. For each summarization paradigm, we first evaluate its sub-summaries separately and then aggregate their scores to derive a paradigm-level score. At the sub-summary level, our framework assesses multiple dimensions. For the two types of background summaries, we evaluate their coverage, focus, and overall quality by comparing them against annotated paragraphs or atomic facts, using paragraph index matching or LLM-based fact verification Wei et al. (2024). For the opinion summary, we assess overall quality by checking its coverage of annotated atomic opinions and identify fine-grained integration errors by classifying failure types for uncovered opinions. We also conduct human evaluation and show a strong correlation with our automatic metrics.
We evaluate 12 advanced LLMs on our benchmark under the structured-prompt Li et al. (2023) and self-reflection Shinn et al. (2023) settings. Our comprehensive results demonstrate that KGDS remains a substantial challenge, with even the top-performing models achieving an average score below 69%. We also identify several key weaknesses: LLMs struggle with the coverage-focus trade-off in background summary retrieval, frequently miss key facts during background summary generation, and fail to resolve implicit references in opinion summary integration. Furthermore, the limited effectiveness of self-reflection indicates that current LLMs lack sufficient self-correction abilities for this task. These findings highlight key bottlenecks and provide concrete directions for future advancements in coarse-grained retrieval, fine-grained generation, and knowledge integration.
@@ 3.1 Preliminary
We establish our benchmark scenario as a two-participant discussion of news content for two reasons. First, news discussions are highly prevalent in daily life, making them more representative than private scenarios such as internal meetings and medical consultations. Second, news summarization Goyal et al. (2022); Zhang et al. (2024); Liu et al. (2024a) is a well-established subfield of automatic summarization research.
@@ 3.2 Annotation
Atomic-Fact-Level Annotation for ABS.  Unlike coarse-grained paragraph annotation, a model-generated ABS is condensed and requires fine-grained, key-point-focused labeling. Inspired by the minimal granularity of atomic facts Liu et al. (2023b); Tang et al. (2024a), we argue that an ideal ABS should cover all key supporting facts while filtering out nonsupporting ones. Motivated by this, we create these two distinct atomic fact sets.
First, we utilize GPT-4o [^3] Version: gpt-4o-2024-08-06 to decompose supporting paragraphs into candidate atomic facts based on principles of indivisibility, independence, and declarativity (see Appendix F.2). Two experts then independently classify each fact as either key or non-key. Through a consistency check, only facts unanimously labeled as key by both experts are considered as the gold standard, resulting in 1,638 key supporting atomic facts.
@@
Second, we use the same decomposition method to nonsupporting paragraphs to generate an initial set of nonsupporting facts. From this set, we then mask [^4] The masking process is consistent with automatic fact verification Tang et al. (2024a), and the masked conflicting facts are not considered in ABS evaluation. any conflicting facts, defined as those that could also be inferred from supporting paragraphs. Such conflicts typically arise when identical or similar facts appear in both supporting and nonsupporting paragraphs at the atomic granularity level. For instance, two events have the same timestamp, but one is the background event while the other is not. By masking these facts, we ensure that the key supporting and nonsupporting fact sets remain non-overlapping, thereby avoiding external verification issues during evaluation. This process yields 4,996 nonsupporting atomic facts, with conflicting facts proving to be sparse at just 176 instances (3.39%).
@@ 4 Evaluation Framework
Our framework is comprehensive and hierarchical. For each of the two summarization paradigms (§2), we independently evaluate the sub-summaries and then aggregate their quality to assess overall performance at the paradigm level. The evaluation methods are interpretable, and the metrics are fine-grained. Figure 2 shows an example. Below, we first describe the evaluation dimensions [^5] We do not focus on fluency and coherence because current LLMs perform well on these dimensions Song et al. (2025). at the sub-summary and paradigm levels (§4.1), and then introduce our automatic methods and metrics (§4.2).
@@ 4.2 Methods and Metrics
ABS Evaluation.  We employ an LLM-based verifier Wei et al. (2024) to check if the summary entails our annotated atomic facts. This process determines whether the summary includes key supporting facts while excluding nonsupporting ones. We use three fine-grained metrics, Key Supporting Atomic Fact Recall, Precision, and F1-score, to quantify the three dimensions of the summary. We provide detailed formulas in Appendix C.1.
AOS Evaluation.  Similar to ABS, we utilize an opinion verifier to measure its coverage of our annotated atomic opinions. We utilize Clear Atomic Opinion Recall to assess its overall quality. For each uncovered atomic opinion, we perform LLM-based error detection to identify the specific integration failure. Specifically, we define five fine-grained error types: implicit reference unclarified, implicit reference incorrectly clarified, opinion misattribution, opinion fact inconsistency, and opinion sentiment distortion. Detailed formulas and error definitions are provided in Appendix C.2 and F.6.
@@ 5 Experimental Setup
We benchmark LLMs for KGDS under two interaction patterns: (1) Single-turn Structured-prompt: A well-structured standard prompt Li et al. (2023) that includes input content, input definition, task description, output definition, and return format (Appendix F.3). (2) Multi-turn Self-reflection: A second-round self-reflection instruction Shinn et al. (2023) with step-by-step chain-of-thought reasoning Wei et al. (2022) following the structured prompt (Appendix F.4).
@@
We use GPT-4o [^8] Version: gpt-4o-2024-11-20 to perform fact and opinion verification as well as error detection due to its excellent consistency with human judgment Song et al. (2024a, b). All prompts are provided in Appendices F.5 and F.6.
@@ 6.1 Background Summary
From Table 3, we find that most LLMs achieve moderate retrieval performance ($\mathrm{SP}_{\mathrm{F_{1}}}\in[71,82]$) and lightweight models (i.e., GPT-4o-mini and Claude 3.5 Haiku) perform poorly. However, from $\mathrm{SP}_{\mathrm{R}}$ and $\mathrm{SP}_{\mathrm{P}}$, we observe significant polarization among LLMs, which indicates distinct retrieval strategies: some prioritize precision at the cost of recall (e.g., GPT-4-turbo), while others do the exact opposite (e.g., DeepSeek-V3). Such imbalance reveals that the current LLMs struggle with the precision-recall trade-off. Moreover, we investigate the paragraph retrieval ratio [^9] The ratio of the number of retrieved paragraphs to the total number of paragraphs. (Figure 3) and identify that most LLMs exhibit either under- or over-retrieval, which is consistent with imbalance.
@@ 6.3 Summarization Paradigm
To validate this, we conduct a human evaluation of overall paradigm-level quality [^10] We focus on paradigm-level validation because the KGDS task’s success depends on the complementarity of both sub-summaries. Furthermore, the individual sub-summary evaluations are already grounded in expert annotations and interpretable, fine-grained verification methods that are known to correlate well with human assessment Song et al. (2024a).. We randomly sample 120 summary pairs (10 per model and 5 per paradigm) from all outputs, and two human evaluators rate their overall quality on a 1-5 Likert scale Joshi et al. (2015), using the average as the final human score. For comparison, we select three LLMs (i.e., GPT-4o, Llama-3.1-405B, DeepSeek-V3) as LLM-as-a-judge baselines to directly score the summaries, with prompt settings following G-Eval Liu et al. (2023a). Table 4 presents the Pearson, Spearman, and Kendall correlation coefficients. Our framework achieves significantly higher correlation with human judgments than the baselines, demonstrating the effectiveness of our fine-grained, hierarchical approach.
@@ 6.4 Self-Reflection Impact
[…]
@@
[…]
@@ 7 Related Work
Dialogue Summarization.  This task spans diverse scenarios, including online chats Gliwa et al. (2019), meetings Hu et al. (2023), and customer service Lin et al. (2021). Existing methods, whether based on feature modeling Chen and Yang (2021); Fang et al. (2022); Lin et al. (2022), pre-training Zou et al. (2021); Zhong et al. (2022), or LLMs Wang et al. (2023a); Zhu et al. (2025), focus heavily on single-source dialogue inputs while neglecting shared background knowledge among participants. Moreover, current evaluation systems Ramprasad et al. (2024); Tang et al. (2024b) also assume single-source inputs. In contrast, our KGDS focuses on multi-source inputs and complementary outputs.
Knowledge-Intensive Dialogue Response.  This task aims to enhance the dialogue process by incorporating external knowledge sources Lewis et al. (2020); Chen et al. (2024). It differs from our KGDS in two key aspects. First, knowledge-intensive dialogue focuses on improving the conversational experience for participants Wang et al. (2023b, 2024), whereas KGDS is designed to generate clear summaries for external readers. Second, the former addresses knowledge gaps between participants, making the dialogue inherently knowledge-intensive Zhang et al. (2020). In contrast, KGDS centers on discussions among participants with shared knowledge, resulting in background-sparse conversations that rely heavily on implicit context.
Summarization Evaluation.  Earlier similarity-based metrics like Rouge Lin (2004), BERTScore Zhang et al. (2019), and MoverScore Zhao et al. (2019) are often misaligned with human judgment. Recently, LLM-based evaluators Chen et al. (2023); Shen et al. (2023); Fu et al. (2024) have been used, yet they still lack interpretability. To address this, some studies Song et al. (2024a); Lee et al. (2024); Yang et al. (2024); Scirè et al. (2024) utilize atomic facts Min et al. (2023) and LLM-based claim verification Song et al. (2024b), achieving more fine-grained evaluation. Among these, FineSurE Song et al. (2024a) is most relevant to our work. However, its metrics lack consistency in evaluation granularity. In contrast, our coverage and focus metrics are both measured at the atomic fact level, ensuring consistent granularity across dimensions.
@@ 8 Conclusion
In this study, we introduce the KGDS task, which aims to create observer-centric summaries by integrating shared background knowledge with discussions. We establish the first benchmark for KGDS and propose a novel hierarchical evaluation framework with fine-grained, interpretable metrics. Our evaluation of 12 leading LLMs reveals significant challenges in background summary retrieval, generation, and opinion summary integration, with even the most advanced models achieving less than 69% average performance across both paradigms.
@@
Our current work is focused on English, creating a monolingual benchmark that may not fully address the challenges of multilingual KGDS. This limitation could impact model performance in languages with different linguistic structures or sociocultural contexts. Nevertheless, English is the most prevalent language in both academic research and real-world LLM applications, making it a reasonable starting point. We plan to explore multilingual and cross-lingual KGDS in future research.
@@
Our KGDS benchmark is developed with careful consideration of ethical implications. The news articles used as shared background knowledge are publicly accessible through Google News and have been carefully reviewed to ensure they do not contain sensitive or harmful information. Explicit consent was obtained from all expert participants involved in discussion construction and data annotation, with a clear explanation of the research purpose and data usage protocols. Our evaluation methodology prioritizes factual accuracy and opinion fidelity to minimize the risk of hallucination propagation in real-world applications. Potential misuse risks, such as generating misleading summaries through improper background-discussion combinations, are mitigated through technical safeguards in our released code and explicit usage guidelines. Researchers using our KGDS benchmark should adhere to responsible AI principles, especially when applying similar techniques to sensitive domains like healthcare or legal discussions. We provide the annotation and evaluation costs as:
@@
Atomic Fact Decomposition 212121
Conflicting Fact Masking 161616
Structured Prompt 747474
Reflection Instruction 126126126
Atomic Fact Verification 173173173
Atomic Opinion Verification 158158158
Fine-Grained Error Detection 898989
Total 657657657
@@
Annotation Cost for KGDS BenchMark.
Our annotation process is multi-step, fine-grained, and highly complex. For each sample in our benchmark, the average annotation time is 2.62.62.6 hours (including news reading and understanding, discussion construction, supporting paragraphs, key supporting atomic facts, nonsupporting atomic facts, and clear atomic opinions annotation). We pay each annotation expert a wage of \$999 per hour (above the minimum wage standard), and the total annotation cost for the KGDS benchmark is \$468046804680 (2.62.62.6 hours per sample $\times$ \$999 per hour $\times$ 222 experts per sample $\times$ 100100100 samples $=$ \$468046804680). The evaluation cost of API calls is detailed in Table 6.

### v3

@@ 1 Introduction
Dialogue summarization aims to distill key topics and interactions from a dialogue into a concise summary kirstein2024cads; rennard2023abstractive; jia2023taxonomy. Conventional paradigm relies primarily on dialogue content, whether in benchmark construction gliwa-etal-2019-samsum; chen-etal-2021-dialogsum; zhu-etal-2021-mediasum, methodologies zhou-etal-2023-multi; tian-etal-2024-dialogue; lu2025mutual; zhu-etal-2025-factual, or evaluation wang-etal-2022-analyzing; gao-wan-2022-dialsummeval; zhu-etal-2023-annotating; tang-etal-2023-context; tang2024tofueval; liu2024exploring; ramprasad2024analyzing. These efforts implicitly assume that "the dialogue itself contains sufficient information to generate a clearly understandable summary for readers."
@@
To advance research in this field, we construct the first KGDS benchmark, situated in a common and realistic scenario of news discussions. It consists of event-rich news articles, each paired with a human-authored discussion. To enable robust assessment, we develop multi-granularity gold evaluation components, which are annotated by experts under strict consistency controls. For the extractive background summary, we annotate coarse-grained supporting and nonsupporting paragraphs. For the abstractive background summary, we create two sets of finest-grained atomic facts min-etal-2023-factscore, including key supporting and nonsupporting facts. For the opinion summary, we introduce clear atomic opinions as basic evaluation units that are minimized and have implicit references clarified.
Furthermore, we propose a novel hierarchical framework for evaluation. For each summarization paradigm, we first evaluate its sub-summaries separately and then aggregate their scores to derive a paradigm-level score. At the sub-summary level, our framework assesses multiple dimensions. For the two types of background summaries, we evaluate their coverage, focus, and overall quality by comparing them against annotated paragraphs or atomic facts, using paragraph index matching or LLM-based fact verification wei2024longformfactualitylargelanguage. For the opinion summary, we assess overall quality by checking its coverage of annotated atomic opinions and identify fine-grained integration errors by classifying failure types for uncovered opinions. We also conduct human evaluation and show a strong correlation with our automatic metrics.
We evaluate 12 advanced LLMs on our benchmark under the structured-prompt li2023structuredchainofthoughtpromptingcode and self-reflection NEURIPS2023_1b44b878 settings. Our comprehensive results demonstrate that KGDS remains a substantial challenge, with even the top-performing models achieving an average score below 69%. We also identify several key weaknesses: LLMs struggle with the coverage-focus trade-off in background summary retrieval, frequently miss key facts during background summary generation, and fail to resolve implicit references in opinion summary integration. Furthermore, the limited effectiveness of self-reflection indicates that current LLMs lack sufficient self-correction abilities for this task. These findings highlight key bottlenecks and provide concrete directions for future advancements in coarse-grained retrieval, fine-grained generation, and knowledge integration.
@@ 3.1 Preliminary
We establish our benchmark scenario as a two-participant discussion of news content for two reasons. First, news discussions are highly prevalent in daily life, making them more representative than private scenarios such as internal meetings and medical consultations. Second, news summarization goyal2022news; zhang2024benchmarking; liu-etal-2024-benchmarking is a well-established subfield of automatic summarization research.
@@ 3.2 Annotation
Atomic-Fact-Level Annotation for ABS.  Unlike coarse-grained paragraph annotation, a model-generated ABS is condensed and requires fine-grained, key-point-focused labeling. Inspired by the minimal granularity of atomic facts liu-etal-2023-revisiting; tang-etal-2024-minicheck, we argue that an ideal ABS should cover all key supporting facts while filtering out nonsupporting ones. Motivated by this, we create these two distinct atomic fact sets.
First, we utilize GPT-4o333gpt-4o-2024-08-06 to decompose supporting paragraphs into candidate atomic facts based on principles of indivisibility, independence, and declarativity (see Appendix G.2). Two experts then independently classify each fact as either key or non-key. Through a consistency check, only facts unanimously labeled as key by both experts are considered as the gold standard, resulting in 1,638 key supporting atomic facts.
@@
Second, we use the same decomposition method to nonsupporting paragraphs to generate an initial set of nonsupporting facts. From this set, we then mask [^4] The masking process is consistent with automatic fact verification tang-etal-2024-minicheck, and the masked conflicting facts are not considered in ABS evaluation. any conflicting facts, defined as those that could also be inferred from supporting paragraphs. Such conflicts typically arise when identical or similar facts appear in both supporting and nonsupporting paragraphs at the atomic granularity level. For instance, two events have the same timestamp, but one is the background event while the other is not. By masking these facts, we ensure that the key supporting and nonsupporting fact sets remain non-overlapping, thereby avoiding external verification issues during evaluation. This process yields 4,996 nonsupporting atomic facts, with conflicting facts proving to be sparse at just 176 instances (3.39%).
@@ 4 Evaluation Framework
Our framework is comprehensive and hierarchical. For each of the two summarization paradigms (§2), we independently evaluate the sub-summaries and then aggregate their quality to assess overall performance at the paradigm level. The evaluation methods are interpretable, and the metrics are fine-grained. Figure 2 shows an example. Below, we first describe the evaluation dimensions [^5] We do not focus on fluency and coherence because current LLMs perform well on these dimensions song2025learningsummarizellmgeneratedfeedback. at the sub-summary and paradigm levels (§4.1), and then introduce our automatic methods and metrics (§4.2).
@@ 4.2 Methods and Metrics
ABS Evaluation.  We employ an LLM-based verifier wei2024longformfactualitylargelanguage to check if the summary entails our annotated atomic facts. This process determines whether the summary includes key supporting facts while excluding nonsupporting ones. We use three fine-grained metrics, Key Supporting Atomic Fact Recall, Precision, and F1-score, to quantify the three dimensions of the summary. We provide detailed formulas in Appendix C.1.
AOS Evaluation.  Similar to ABS, we utilize an opinion verifier to measure its coverage of our annotated atomic opinions. We utilize Clear Atomic Opinion Recall to assess its overall quality. For each uncovered atomic opinion, we perform LLM-based error detection to identify the specific integration failure. Specifically, we define five fine-grained error types: implicit reference unclarified, implicit reference incorrectly clarified, opinion misattribution, opinion fact inconsistency, and opinion sentiment distortion. Detailed formulas and error definitions are provided in Appendix C.2 and G.6.
@@ 5 Experimental Setup
We benchmark LLMs for KGDS under two interaction patterns: (1) Single-turn Structured-prompt: A well-structured standard prompt li2023structuredchainofthoughtpromptingcode that includes input content, input definition, task description, output definition, and return format (Appendix G.3). (2) Multi-turn Self-reflection: A second-round self-reflection instruction NEURIPS2023_1b44b878 with step-by-step chain-of-thought reasoning NEURIPS2022_9d560961 following the structured prompt (Appendix G.4).
@@
We use GPT-4o888gpt-4o-2024-11-20 to perform fact and opinion verification as well as error detection due to its excellent consistency with human judgment song-etal-2024-finesure; song-etal-2024-veriscore. All prompts are provided in Appendices G.5 and G.6.
@@ 6.1 Background Summary
From Table 3, we find that most LLMs achieve moderate retrieval performance ($\mathrm{SP}_{\mathrm{F_{1}}}\in[71,82]$) and lightweight models (i.e., GPT-4o-mini and Claude 3.5 Haiku) perform poorly. However, from $\mathrm{SP}_{\mathrm{R}}$ and $\mathrm{SP}_{\mathrm{P}}$, we observe significant polarization among LLMs, which indicates distinct retrieval strategies: some prioritize precision at the cost of recall (e.g., GPT-4-turbo), while others do the exact opposite (e.g., DeepSeek-V3). Such imbalance reveals that the current LLMs struggle with the precision-recall trade-off. Moreover, we investigate the paragraph retrieval ratio [^9] Defined as the ratio of the number of retrieved paragraphs to the total number of paragraphs. (Figure 3) and identify that most LLMs exhibit either under- or over-retrieval, which is consistent with imbalance.
@@ 6.3 Summarization Paradigm
To validate this, we conduct a human evaluation of overall paradigm-level quality [^10] We focus on paradigm-level validation because the KGDS task’s success depends on the complementarity of both sub-summaries. Furthermore, the individual sub-summary evaluations are already grounded in expert annotations and interpretable, fine-grained verification methods that are known to correlate well with human assessment song-etal-2024-finesure.. We randomly sample 120 summary pairs (10 per model and 5 per paradigm) from all outputs, and two human evaluators rate their overall quality on a 1-5 Likert scale joshi2015likert, using the average as the final human score. For comparison, we select three LLMs (i.e., GPT-4o, Llama-3.1-405B, DeepSeek-V3) as LLM-as-a-judge baselines to directly score the summaries, with prompt settings following G-Eval liu-etal-2023-g. Table 4 presents the Pearson, Spearman, and Kendall correlation coefficients. Our framework achieves significantly higher correlation with human judgments than the baselines, demonstrating the effectiveness of our fine-grained, hierarchical approach.
@@ 6.4 Self-Reflection Impact
From Table 3, we observe that the performance fluctuations (i.e., $\uparrow$} or $\downarrow$}) after self-reflection are limited (the maximum average fluctuation across all metrics is $2.30$ for $\mathrm{CAO}_{\mathrm{R\text{:ABS-AOS}}}$). This means that self-reflection does not fundamentally influence the capacities of LLMs in KGDS, revealing two key limitations: (1) LLMs lack sufficient self-evaluation abilities for KGDS. (2) The reflection strategies of LLMs struggle to provide excellent reasoning paths for KGDS.
@@
As presented in Table 5, most LLMs exhibit a decrease in IRIC and an increase in IRU (e.g., GPT-4o with IRIC $\downarrow 2.09$}, IRU $\uparrow 2.80$}). This suggests that self-reflection encourages LLMs to adopt more cautious strategies for clarifying implicit references — opting to leave them unclarified rather than risk incorrect clarification. However, such conservatism ultimately harms the quality of opinion summaries, as evidenced by the widespread drop in $\mathrm{CAO}_{\mathrm{R}}$ in Table 3.
@@ 7 Related Work
Dialogue Summarization.  This task spans diverse scenarios, including online chats gliwa-etal-2019-samsum, meetings hu-etal-2023-meetingbank, and customer service lin-etal-2021-csds. Existing methods, whether based on feature modeling chen-yang-2021-structure; fang-etal-2022-spoken; lin-etal-2022-roles, pre-training zou-etal-2021-low; zhong2022dialoglm, or LLMs wang-etal-2023-instructive; zhu-etal-2025-factual, focus heavily on single-source dialogue inputs while neglecting shared background knowledge among participants. Moreover, current evaluation systems ramprasad2024analyzing; tang2024tofueval also assume single-source inputs. In contrast, our KGDS focuses on multi-source inputs and complementary outputs.
Knowledge-Intensive Dialogue Response.  This task aims to enhance the dialogue process by incorporating external knowledge sources NEURIPS2020_6b493230; chen2024benchmarking. It differs from our KGDS in two key aspects. First, knowledge-intensive dialogue focuses on improving the conversational experience for participants wang-etal-2023-large; wang2024improving, whereas KGDS is designed to generate clear summaries for external readers. Second, the former addresses knowledge gaps between participants, making the dialogue inherently knowledge-intensive zhang-etal-2020-grounded. In contrast, KGDS centers on discussions among participants with shared knowledge, resulting in background-sparse conversations that rely heavily on implicit context.
Summarization Evaluation.  Earlier similarity-based metrics like Rouge lin2004rouge, BERTScore zhang2019bertscore, and MoverScore zhao-etal-2019-moverscore are often misaligned with human judgment. Recently, LLM-based evaluators chen-etal-2023-exploring-use; shen-etal-2023-large; fu-etal-2024-gptscore have been used, yet they still lack interpretability. To address this, some studies song-etal-2024-finesure; lee-etal-2024-unisumeval; yang-etal-2024-fizz; scire-etal-2024-fenice utilize atomic facts min-etal-2023-factscore and LLM-based claim verification song-etal-2024-veriscore, achieving more fine-grained evaluation. Among these, FineSurE song-etal-2024-finesure is most relevant to our work. However, its metrics lack consistency in evaluation granularity. In contrast, our coverage and focus metrics are both measured at the atomic fact level, ensuring consistent granularity across dimensions.
@@ 8 Conclusion
In this study, we introduce the KGDS task, which aims to create observer-centric summaries by integrating shared background knowledge with discussions. We establish the first benchmark for KGDS and propose a task-aligned hierarchical evaluation framework with fine-grained, interpretable metrics. Our evaluation of 12 leading LLMs reveals significant challenges in background summary retrieval, generation, and opinion summary integration, with even the most advanced models achieving less than 69% average performance across both paradigms.
@@
Our study presents several limitations that merit discussion and future exploration:
Domain Generalization.
Knowledge-grounded discussions and their confusing summaries are prevalent in both open-domain and private scenarios. Our benchmark is centered on open-domain news discussions, which may limit its applicability to private scenarios (e.g., internal meetings, medical consultations, legal debates, etc.). However, shared background knowledge and the related discussions in private contexts are challenging to obtain and are less common and representative than our open benchmark. Future work will focus on validating the performance of LLMs in handling KGDS across diverse private scenarios.
Language Diversity.
Our current research is focused on English, creating a monolingual benchmark that may not fully address the challenges of multilingual KGDS. This limitation could impact model performance in languages with different linguistic structures or sociocultural contexts. Nevertheless, English is the most prevalent language in both academic research and real-world LLM applications, making it a reasonable starting point. We plan to explore multilingual and cross-lingual KGDS in future research.
@@
Our KGDS benchmark is developed with careful consideration of ethical implications. The news articles used as shared background knowledge are publicly accessible through Google News and have been carefully reviewed to ensure they do not contain sensitive or harmful information. Explicit consent was obtained from all expert participants involved in discussion construction and data annotation, with a clear explanation of the research purpose and data usage protocols. Our evaluation methodology prioritizes factual accuracy and opinion fidelity to minimize the risk of hallucination propagation in real-world applications. Potential misuse risks, such as generating misleading summaries through improper background-discussion combinations, are mitigated through technical safeguards in our released code and explicit usage guidelines. Researchers using our KGDS benchmark should adhere to responsible AI principles, especially when applying similar techniques to sensitive domains like healthcare or legal discussions.
Annotation and Evaluation Cost.
Our annotation process is multi-step, fine-grained, and highly complex. For each sample in our benchmark, the average annotation time is $2.6$ hours (including news reading and understanding, discussion construction, supporting paragraphs, key supporting atomic facts, nonsupporting atomic facts, and clear atomic opinions annotation). We pay each annotation expert a wage of \$99 per hour (above the minimum wage standard), and the total annotation cost for the KGDS benchmark is \$$4680$ ($2.6$ hours per sample $\times$ \$99 per hour $\times$ 22 experts per sample $\times$ $100$ samples == \$$4680$). The evaluation cost of API calls is detailed in Table 6.
@@
Atomic Fact Decomposition 2121
Conflicting Fact Masking $16$
Structured Prompt $74$
Reflection Instruction $126$
Atomic Fact Verification $173$
Atomic Opinion Verification $158$
Fine-Grained Error Detection $89$
Total $657$